Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-12x + 27y = 60
-6x – 6y = 30

Answer by mananth(16949)

-12 x + 27 y = 60 .............1
-6 x -6 y = 30 .............2
Eliminate y
multiply (1)by 6
Multiply (2) by 27
-72 x 162 y = 360
-162 x + -162 y = 810
Add the two equations
-234 x = 1170.00
/ -234
x = -5
plug value of x in (1)
-12 x + 27 y = 60
60 + 27 y = 60
27 y = 60 -60
27 y = 0
y = 0
m.ananth@hotmail.ca